OpenStudy (tori1423):

Square RSTU is inscribed in circle P. Given the coordinates for the vertices of the square, find the equation of circle P. R(0, 4), S(6, 2), T(4, -4), and U(-2, -2) (x - 2)2 + y2 = 68 (x + 2)2 + y2 = 20 (x - 2)2 + y2 = 20

OpenStudy (campbell_st):

well since the square is inside the circle... find the midpoint of a diagonal... that will be the centre of the circle.. then you can find the equation. to check find the midpoint of the other diagonal... and it should be the same point as the 1st diagonal... the reason is that diagonals in a square bisect each other. hope it helps

OpenStudy (campbell_st):

now one of the vertices of the square is on the circumference of the circle. so substitute a point into your equation to find r^2 then you'll have the correct answer

OpenStudy (tori1423):

K

OpenStudy (tori1423):

2,0 is the midpoint @campbell_st

OpenStudy (tori1423):

i need help finding radius though

OpenStudy (campbell_st):

great so the general equation is \[(x - h)^2 + (y - k)^2 = r^2\] where (h, k) is the centre so using your midpoint the equation is \[(x -2)^2 + (y - 0)^2 = r^2\] to find the value of r^2 substitute a point that is a vertex... as it's on the circumference then you'll have all the bits for the equation it doesn't matter which vertex you choose

OpenStudy (tori1423):

ok

OpenStudy (tori1423):

so 20 @campbell_st

OpenStudy (campbell_st):

correct... so do you have an answer choice that matches... ?

OpenStudy (tori1423):

C

OpenStudy (campbell_st):

CC

OpenStudy (tori1423):

thxs